The age of Betty's mother is 6 years more than four times Betty's age. The difference in their ages is 21 years. How old is Betty and her mother? Show the linear equation in one variable used to solve it and then show and explain each step in solving that equation for your answer.
Answer by stanbon(75887)

The age of Betty's mother is 6 years more than four times Betty's age. The difference in their ages is 21 years. How old is Betty and her mother?
-----
Equations:
m = 4b+6
m - b = 21
-----
Substitute for "b" so you can solve for "m":
4b + 6 = b + 21
3b = 15
b = 5 yrs (Betty's age)
------
Solve for "m":
m = b+21
m = 26 yrs (Mother's age)
==========================
Cheers,
Stan H.